### v1

# A fractal geometry immersed in a hierarchical magnetic flux distribution

Abstract
Fractal geometry presents us with a self-similarity in their pattern at various length scales that is prevalent in our natural world. We present theoretical model of a Sierpinski gasket (SPG) fractal geometry with a deterministic perturbation in the form of a hierarchical distribution of magnetic flux. Such flux configuration induces a deterministic disorder in the Aharonov-Bohm (AB) phases picked up by the electron wavefunction. Using the tight-binding formalism, we show that by tunning the strength of the hierarchy parameter of those AB phases, one can systematically engineer quantum states in a SPG fractal lattice. In addition to this, we have also observed that by controlling the strength of this hierarchy parameter in the magnetic flux, one can effectively regulate the persistent current in the SPG fractal structure. This characteristic is found to be true for various filling factors. Our results could be useful for designing nanoelectronic devices using molecular fractal structures fabricated by chemical synthesis technique.
I Introduction
The terminology of fractal geometries which was initially emerged purely as a mathematical concept mandelbrot-book , has generated intense research interest in various domains of condensed matter physics and material science in recent times. Although deterministic fractal geometries belongs to the class of aperiodic systems, they are not completely random; they possess certain self-similar beautiful pattern and symmetrical structure in their construction. This often helps theoreticians to analytically investigate various interesting properties appearing in such fractal structures. Also, deterministic growth rule allows one to grow such fractal structures in the laboratory using the state-of-the-art self-assembly or lithography techniques self-assembly-1 ; self-assembly-2 ; hexagonal-SPG , and perform fascinating experiments with them.
One of the key reasons why fractal lattice structures have drawn a lot of research attention of late is due to the fact that they possess non-integer Hausdorff dimension $D_{f}$. This has ignited a lot of curiosity in the mind of the researchers and led to a number of intriguing investigations in this direction, viz., topological character of electronic states on fractals titus ; shreya , flat bands in fractal-like lattices biplab ; atanu ; hahafi-fractal-FB ; chen-fractal-FB-1 ; chen-fractal-FB-2 ; sougata-fractal-FB , Hall conductivity askar-Hall ; cm-smith-Hall and possible existence of anyons in fractal structures biplab-PRR-2020 , are to name a few of them. The rapid advancement in experimental tools and fabrication methods has also played a major role in developing renewed interest in the physics of fractal geometries. These technological progress has enabled us to synthesize tailor-made fractal structures in the laboratory using various techniques, such as, self-assembled molecular growth of fractal materials self-assembly-1 ; self-assembly-2 , designing quantum fractal using atomic manipulation in a scanning tunnelling microscope cm-smith-nat-phys , fractal photonic lattice consisting of waveguides realized using femtosecond-laser-writing technology m-segev etc.
Inspired by these experimental advancements in creating the fractal structures in reality, various interesting theoretical studies concerning the deterministic fractal structures have taken place in recent times. This includes works on energy level statistics in fractal clusters yorikawa-JPC2019 ; askar-PRB2019 , confinement of electrons in fractal plane using electric field yuan-PRB2020 , electronic fractal pattern in building molecular fractal systems latge-pccp2022 , role of anisotropy on magneto-transport in fractal network maiti-PLA2020 , and Josephson effect in a fractal geometry vladimir-arXiv2024 among others.
Figure 1: Schematic diagram of a Sierpinski gasket fractal lattice with hierarchical distribution of a magnetic flux following the pattern $\Phi_{\ell}=\lambda^{\ell}\Phi$, where $\ell$ represents the generation index.
It is expected that, in future, further advancement in nanotechnology and fabrication techniques will enable us to do more delicate experiments with these fractal lattice models. Such possibility inspires us to undertake this theoretical investigation presented in this article. We consider a Sierpinski gasket (SPG) fractal lattice structure immersed in a deterministic hierarchical distribution of magnetic fluxes. The ramification of a SPG fractal structure is such that, it offers us with down-headed triangular loops of gradually increasing size (See Fig. 1). Following this ramification pattern, we introduce a hierarchical distribution of magnetic fluxes $\Phi_{\ell}=\lambda^{\ell}\Phi$ ($\lambda$ being the hierarchy parameter) through these down-headed triangular closed loops. Such distribution of the magnetic fluxes will be manifested as a distribution of Aharonov-Bohm (AB) phases, whenever an electron will hop from one atomic site to another atomic site along these down-headed triangular closed loops. This provides us with an interesting situation, where we have a deterministic disorder in the distribution of the AB phases in this closed-loop fractal geometry controlled by a hierarchy parameter $\lambda$.
In this work, we investigate the effect of this usual distribution of magnetic fluxes on the energy eigenvalue spectrum of the SPG fractal network as a function of the hierarchy parameter $\lambda$. The presence of the magnetic flux through such closed-loop fractal lattice geometry induces a circulating current in these systems known as persistent current maiti-PRB2010 . It is a pertinent question to ask, may we have some mechanism by which we can control the persistent current in such closed-loop fractal lattice geometries? In this study, we address this question and show that, it is indeed possible to effectively control the magnitude of the persistent current in this SPG fractal lattice by tunning the hierarchy parameter $\lambda$ in the distribution of the magnetic flux. Our model may be suitable for mesoscopic or nanoscopic experiments using fractal structures. In what follows, we present the schematic and mathematical framework of our model in Sec. II. We discuss the effect of the hierarchy parameter on the energy eigenvalue spectrum of the SPG fractal lattice model in Sec. III. The role of the hierarchy parameter in controlling the persistent current in such complex systems is elaborated in Sec. IV. Finally, in Sec. V, we draw our conclusion highlighting the summary of our findings and their utility, and the scope of future study in this direction.
II The model and the mathematical framework
We consider a theoretical model of a SPG fractal lattice which follows certain deterministic ramification pattern in its growth rule as shown in Fig. 1. The number of sites in a $\ell$-th generation system is given by the following formula:
$\mathcal{N}(\ell)=\dfrac{3}{2}\big{(}1+R^{\ell}\big{)},$ (1)
where $R$ is the repetition factor (which is equal to 3 for this particular fractal model) and $\ell$ denotes the generation index. The single-electron states in such a lattice structure can be described by the following Hamiltonian within the tight-binding framework:
[…]
where first term represents the on-site energy and the second term represents the nearest-neighbor hopping integral. $\theta_{\ell}=\dfrac{2\pi\Phi_{\ell}}{3(2^{\ell-1})\Phi_{0}}$ is the Aharonov-Bohm phase A-B-Phase associated with the nearest-neighbor hopping parameter $t_{mn}$, whenever an electron hops along a bond in the down-headed triangular loops pierced by an external magnetic flux $\Phi_{\ell}=\lambda^{\ell}\Phi$, where $\lambda$ is the hierarchy parameter. $\Phi$ is measured in units of fundamental flux quantum $\Phi_{0}=hc/e$. $c_{n}^{\dagger}\,(c_{n})$ denotes the creation (annihilation) operator for an electron at a site $n$, and $\epsilon_{n}$ is the on-site potential for the $n$-th atomic site.
Following the Eq. (2), we can construct the Hamiltonian matrix for this fractal lattice model of different system sizes, which will allow us to extract the information about the energy spectrum as well as for the persistent current of this system as a function of the deterministic disorder distribution in the AB phases. It is to be noted that, because of the self-similar characteristics of a fractal system, once we get some interesting properties in such systems for a small system size, it generally retains those properties even when we go to a large system size. Thus, considering the possibility of experimental realization of this model with such an unusual hierarchical distribution of magnetic fluxes in the real-life laboratory, we consider moderately large system for our theoretical calculations. At first, we study the role of the hierarchy parameter $\lambda$ which controls the hierarchical distribution of the magnetic flux $\Phi_{\ell}$ on the energy eigenvalue spectrum for this model. This is elaborated in the next section.
III Effect of the hierarchy parameter on the energy eigenvalue spectrum
In this section, we discuss the role of the hierarchy parameter $\lambda$ on the energy eigenvalue spectrum of the SPG lattice. We have first explicitly constructed the Hamiltonian matrix for the SPG lattice using Eq. (2). We take the on-site potential $\epsilon_{n}=0$ for all the sites, and consider only nearest-neighbor constant coupling between the sites $t_{mn}=t=1$. The direction of the hierarchical fluxes $\Phi_{\ell}$ (where $\ell=1,2,3,\ldots$) in different down-headed triangular closed plaquettes is taken to be in the out-of-plane direction. Accordingly, the electrons will pick up the AB phases ±$\pm\theta_{\ell}$ (‘$+$’ is for forward hopping and ‘--’ is for backward hopping). Once the Hamiltonian matrices are constructed for different generations of the SPG lattice, one can easily diagonalize them to evaluate the energy eigenvalue spectrum for the lattice structure. Here, we present the results for the 3rd and the 4th generation SPG lattice structure in Fig. 2 and Fig. 3, respectively.
Figure 2: Variation of the energy eigenvalue spectrum for a 3rd generation SPG lattice structure [with $\mathcal{N}(3)=42$] as a function of the hierarchical magnetic flux for different values of the hierarchy parameter, viz., (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
We start our discussion with $\lambda=0$, i. e., in absence of the hierarchy parameter. It is clear that for $\lambda=0$, the hierarchical magnetic flux distribution in all the plaquettes vanishes. Hence, under this condition, we get the energy eigenvalue distribution for the SPG lattice geometry in absence of the magnetic flux. The energy eigenvalue spectrum exhibit highly degenerate, discrete, fractal characteristics [see Fig. 2(a)]. Then, we incorporate a small non-zero value of $\lambda=0.1$. With this small non-zero value of $\lambda$, we instantly observe that, the degeneracy in the energy spectrum is slightly lifted [Fig. 2(b)]. As we gradually tune the strength of the hierarchy parameter $\lambda$ more, the energy eigenvalue spectrum become increasingly non-degenerate [Fig. 2(c) and 2(d)]. Thus, by controlling the strength of the hierarchy parameter $\lambda$, one can systematically control the degeneracy of the energy spectrum in a SPG fractal structure.
Figure 3: Energy eigenvalue distribution for a 4th generation SPG lattice structure [with $\mathcal{N}(4)=123$] as a function of the hierarchical magnetic flux with values of the hierarchy parameter as (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
It is to be noted that, the self-similar aspect of a fractal structure is, in general, reflected in all its properties. Therefore, one can expect that, our observations appearing in Fig. 2 for a 3rd generation SPG fractal structure will qualitatively persist even when we go to a higher generation fractal structure. To confirm this fact, we have also evaluated the energy spectrum for a 4th generation SPG fractal with different values of the hierarchy parameter $\lambda$. This is depicted in Fig. 3(a)-(d). It is clearly evident that, the qualitative structure of the energy spectrum remains same as in Fig. 2(a)-(d), apart from appearance of some additional energy eigenvalues which appear in the spectrum due to increase in the system size. We anticipate the similar feature to be sustained even for larger SPG fractal structures. It is worth to point out that, the ramification of the SPG fractal structure is such that one cannot have such hierarchical distribution of magnetic fluxes in the upward triangular plaquettes (see Fig. 1). In upward triangular plaquettes, one can only have a uniform distribution of the magnetic flux as discussed in Ref. maiti-PRB2010 ; chakrabarti-PRB2005 .
Figure 4: Average density of states (ADOS) for a 3rd generation SPG fractal structure as a function of the energy and hierarchical magnetic flux. (a) is for $\lambda=0$ and (b) is for a non-zero value of $\lambda=0.9$.
Figure 5: Variation of the persistent current ($I_{\textrm{max}}$) as a function of the hierarchy parameter $\lambda$ for a 3rd generation SPG fractal lattice. (a) is for a filling factor $\nu=1/3$, (b) is for $\nu=1/2$, and (c) represents $\nu=2/3$.
To corroborate our results on the energy spectrum of the SPG fractal lattice as a function of the hierarchy parameter, we have calculated the average density of states (ADOS) for a finite generation SPG fractal structure. To compute the ADOS, we have used the Green’s function technique which people uses in general for such fractal lattice models nandy-JPCM2015 ; pal-PRB2012 ; pal-PRA2013 ; pal-EPJB2012 . Using the Green’s function technique, ADOS can be defined as,
$\rho(E,\Phi)=-\dfrac{1}{\mathcal{N}\pi}\textrm{Im}\Big{[}\textrm{Tr}\left[\bm{\mathcal{G}}(E,\Phi)\right]\Big{]},$ (3)
[…]
It is important to see the manifestation of these properties in terms of some experimentally measurable quantities. This brings us to the next section, where we demonstrate how one can control the persistent current as a function of the hierarchy parameter in this closed-loop fractal geometry.
IV Controlling the persistent current
In isolated mesoscopic metallic rings threaded by a magnetic flux, one can have a dissipationless circulating current termed as persistent current buttiker-PLA1983 ; gefen-PRL1989 . It is important to emphasize the fact that, this current has nothing to do with the persistent current found in a superconducting material. This idea was initially proposed for a single mesoscopic conducting loop structure gefen-PRL1989 ; georges-bookchapter2007 , but later on people have also investigated the behavior of such persistent current in multi-loop fractal geometries maiti-PRB2010 . The persistent current as a function of the magnetic flux $\Phi$ is related to the energy eigenvalues (energy levels) $E_{\alpha}$ at absolute zero temperature ($T=0$ K) as follows:
$I_{\alpha}(\Phi)=-c\;\dfrac{\partial E_{\alpha}(\Phi)}{\partial\Phi},$ (4)
where $c$ is a constant which is set equal to 1 for our calculation. The total persistent current in the system is given by,
$I(\Phi)=\sum_{\alpha=1}^{\mathcal{N}_{e}}I_{\alpha}=-\sum_{\alpha=1}^{\mathcal{N}_{e}}\dfrac{\partial E_{\alpha}}{\partial\Phi},$ (5)
where $\mathcal{N}_{e}$ is the total number of electrons in the system. We define $\nu=\mathcal{N}_{e}/\mathcal{N}(\ell)$ as the “filling factor” for a $\ell$-th generation SPG fractal system. Note that, we consider non-interacting spinless electron model.
For our model, the magnetic flux $\Phi$ depends on the hierarchy parameter $\lambda$. So, we can study the magnetic response of the SPG fractal geometry in terms of the persistent current as a function of $\lambda$. We have systematically investigated the behavior of the persistent current in a 3rd generation SPG fractal network as a function of the hierarchy parameter $\lambda$ for different values of $\nu$. This is illustrated in Fig. 5(a)-(c). It is observed that, for different filling factors $\nu$, the persistent current in the system gets enhanced as one increases the value of $\lambda$. This opens up an interesting possibility by which one can flexibly control the persistent current in such multi-loop fractal geometries. Once again, this feature is also tested for a 4th generation SPG fractal structure (Fig. 6), and found to be true as an evidence of the generic self-similar property of a fractal structure.
Figure 6: Variation of the persistent current ($I_{\textrm{max}}$) as a function of the hierarchy parameter $\lambda$ for a 4th generation SPG fractal structure with (a) $\nu=1/3$ and (b) $\nu=2/3$.
It is to be noted that, in principle, one could study these properties for any other higher generation SPG fractal structures. We predict that, our results will be qualitatively retained for those cases also. However, from the experimental point-of-view, one cannot go beyond certain system size in the real-life laboratory when it comes to fabricating such complex fractal geometries.
V Summary and future outlook
In conclusion, we have studied the role of a hierarchical distribution of magnetic flux on the energy spectrum and persistent current in a multi-loop Sierpinski gasket fractal structure. We have shown that, by controlling the hierarchy parameter of the embedded magnetic flux in this SPG fractal network, one can effectively engineer quantum states in such a complex system. From the application point-of-view, we have shown that this directly impacts the behavior of the persistent current in this SPG fractal structure. We can suitably control the persistent current in the system by tunning the hierarchy parameter in a flexible way for different filling factors of the electronic states. One could in principle, test these results by designing an experiment using a self-assembled molecular SPG fractal structure. This could be helpful to devise an unconventional nanoelectronic device using a molecular SPG fractal structure. In future, we plan to generalize this idea for other similar kind of closed-loop fractal geometries, such as Vicsek fractal pal-PRB2012 ; pal-PRA2013 .
Acknowledgements.
The author would like to thank Prof. Arunava Chakrabarti from Presidency University, Kolkata for his useful feedback on the initial stage of this work. The author also acknowledges helpful discussion with Amrita Mukherjee from TIFR, Mumbai on the calculation of the persistent current.

### v2

@@
 Inspired by these experimental advancements in creating the fractal structures in reality, various interesting theoretical studies concerning the deterministic fractal structures have taken place in recent times.
-This includes works on energy level statistics in fractal clusters yorikawa-JPC2019 ; askar-PRB2019 , confinement of electrons in fractal plane using electric field yuan-PRB2020 , electronic fractal pattern in building molecular fractal systems latge-pccp2022 , role of anisotropy on magneto-transport in fractal network maiti-PLA2020 , and Josephson effect in a fractal geometry vladimir-arXiv2024 among others.
+This includes works on energy level statistics in fractal clusters yorikawa-JPC2019 ; askar-PRB2019 , confinement of electrons in fractal plane using electric field yuan-PRB2020 , electronic fractal pattern in building molecular fractal systems latge-pccp2022 , role of anisotropy on magneto-transport in fractal network maiti-PLA2020 , Josephson effect in a fractal geometry vladimir-APL2024 , and stability of loop current states in fractal structures built with Bose-Einstein condensates anna-PRA2024 among others.
 Figure 1: Schematic diagram of a Sierpinski gasket fractal lattice with hierarchical distribution of a magnetic flux following the pattern $\Phi_{\ell}=\lambda^{\ell}\Phi$, where $\ell$ represents the generation index.
@@
 This provides us with an interesting situation, where we have a deterministic disorder in the distribution of the AB phases in this closed-loop fractal geometry controlled by a hierarchy parameter $\lambda$.
+It is to be noted that, in Fig.
+1, although schematically we have shown that the fluxes are embedded in the downward triangles, but in reality, along the arms of the upward triangles, we will also have various distribution of the Aharonov-Bohm phases depending on their position in the fractal geometry for any given finite generation.
+This means that, the upward triangles are also containing some effective fluxes.
+One may realize such a complex distribution of magnetic fluxes by putting up some synthetic gauge fields gardin-arXiv2024 in various downward plaquettes, which may be controlled locally in each plaquette.
 In this work, we investigate the effect of this usual distribution of magnetic fluxes on the energy eigenvalue spectrum of the SPG fractal network as a function of the hierarchy parameter $\lambda$.
@@
 3, respectively.
-Figure 2: Variation of the energy eigenvalue spectrum for a 3rd generation SPG lattice structure [with $\mathcal{N}(3)=42$] as a function of the hierarchical magnetic flux for different values of the hierarchy parameter, viz., (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
+We note that, the energy ($E$) of the particle is measured in units of the hopping integral $t$.
+Figure 2: Variation of the energy eigenvalue spectrum for a 3rd generation SPG lattice structure (with $\mathcal{N}(3)=42$) as a function of the hierarchical magnetic flux for different values of the hierarchy parameter, viz., (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
 We start our discussion with $\lambda=0$, i.
@@
 Thus, by controlling the strength of the hierarchy parameter $\lambda$, one can systematically control the degeneracy of the energy spectrum in a SPG fractal structure.
-Figure 3: Energy eigenvalue distribution for a 4th generation SPG lattice structure [with $\mathcal{N}(4)=123$] as a function of the hierarchical magnetic flux with values of the hierarchy parameter as (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
+Figure 3: Energy eigenvalue distribution for a 4th generation SPG lattice structure (with $\mathcal{N}(4)=123$) as a function of the hierarchical magnetic flux with values of the hierarchy parameter as (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
 It is to be noted that, the self-similar aspect of a fractal structure is, in general, reflected in all its properties.
@@
 (a) is for $\lambda=0$ and (b) is for a non-zero value of $\lambda=0.9$.
-Figure 5: Variation of the persistent current ($I_{\textrm{max}}$) as a function of the hierarchy parameter $\lambda$ for a 3rd generation SPG fractal lattice.
+Figure 5: Variation of the persistent current ($I_{\textrm{max}}$) as a function of the hierarchy parameter $\lambda$ for a 3rd generation SPG fractal lattice for the temperature $T=0$ and $T>0$, respectively.
 (a) is for a filling factor $\nu=1/3$, (b) is for $\nu=1/2$, and (c) represents $\nu=2/3$.
@@
 […]
-It is important to see the manifestation of these properties in terms of some experimentally measurable quantities.
+It is important to see the manifestation of the above properties in terms of some experimentally measurable quantities.
 This brings us to the next section, where we demonstrate how one can control the persistent current as a function of the hierarchy parameter in this closed-loop fractal geometry.
@@
 $I_{\alpha}(\Phi)=-c\;\dfrac{\partial E_{\alpha}(\Phi)}{\partial\Phi},$ (4)
-where $c$ is a constant which is set equal to 1 for our calculation. The total persistent current in the system is given by,
+where $c$ is a constant which is set equal to 1 for our calculation.
+We note that, for our model, we have the hierarchical distribution of the magnetic fluxes as $\Phi_{\ell}=\lambda^{\ell}\Phi$, where $\lambda$ is the hierarchy parameter, $\ell$ is the generation index, and $\Phi$ is some fundamental flux which can take values from $-\Phi_{0}$ to $\Phi_{0}$.
+Therefore, $\Phi$ in Eq.
+(4), for our model can be expressed in terms of various orders of $\Phi_{\ell}$ and $\lambda^{\ell}$ depending on any given ‘$\ell$’-th generation fractal system.
+The total persistent current in the system is given by,
 $I(\Phi)=\sum_{\alpha=1}^{\mathcal{N}_{e}}I_{\alpha}=-\sum_{\alpha=1}^{\mathcal{N}_{e}}\dfrac{\partial E_{\alpha}}{\partial\Phi},$ (5)
@@
 6), and found to be true as an evidence of the generic self-similar property of a fractal structure.
+We remark that, in a real experimental setup, if one typically chooses the system size to be $\sim\mu$m, magnetic field $\sim 10$ Tesla, then the persistent current we get will be $\sim$ nano-Amp (nA).
 Figure 6: Variation of the persistent current ($I_{\textrm{max}}$) as a function of the hierarchy parameter $\lambda$ for a 4th generation SPG fractal structure with (a) $\nu=1/3$ and (b) $\nu=2/3$.
+It is important to briefly mention the effect of the finite temperature on the persistent current for our fractal lattice system.
+At finite temperature ($T>0$), one can calculate the persistent current in the system using the following formula bloch-PRB1970 ; wendler-PSSB1995 :
+$I(\Phi)=-\dfrac{\partial F(\Phi)}{\partial\Phi}$ (6a)
+with $F(\Phi)=-k_{B}T\>\ln\Big{(}\sum_{\nu}-\dfrac{E_{\nu}}{k_{B}T}\Big{)}$ (6b)
+where $F$ is the Helmholtz free energy, $k_{B}$ is the Boltzmann constant, and $T$ is the absolute temperature.
+It is found that, at finite temperature, the persistent current in the system decreases drastically as evident from Fig.
+5 and Fig.
+6.
+This is because, as we go to the finite temperature, electrons can occupy the higher energy levels with different phases, which tend to cancel out the net current in the system.
+Figure 7: Variation of the persistent current ($I$) as a function of the magnetic flux for three different values of the hierarchy parameter $\lambda=0.3$ (red solid line), $\lambda=0.6$ (blue dashed line), and $\lambda=0.9$ (green dotted line), respectively in a $3$rd generation SPG fractal structure.
+The filling factor is set be $\nu=1/3$.
+The variation of the persistent current with the magnetic flux for different fixed values of the hierarchy parameter $\lambda$ is shown in Fig.
+7.
+From these plots, we can clearly identify that as the value of $\lambda$ increases, the magnitude of the persistent current in the system also increases.
+This result is consistent with the result shown in Fig.
+5.
+We note that, the periodicity of the persistent current as a function of the magnetic flux also increases as we increase the value of $\lambda$.
+This is because of the hierarchical nature of the magnetic flux distribution in the system as a function of the hierarchy parameter $\lambda$.
 It is to be noted that, in principle, one could study these properties for any other higher generation SPG fractal structures.
@@
 In future, we plan to generalize this idea for other similar kind of closed-loop fractal geometries, such as Vicsek fractal pal-PRB2012 ; pal-PRA2013 .
+Another interesting future prospect of this work could be to study these properties in fractal lattices with self-affine feature self-affine .
 Acknowledgements.
@@
 Arunava Chakrabarti from Presidency University, Kolkata for his useful feedback on the initial stage of this work.
-The author also acknowledges helpful discussion with Amrita Mukherjee from TIFR, Mumbai on the calculation of the persistent current.
+The author also acknowledges helpful discussion with Dr.
+Amrita Mukherjee from TIFR, Mumbai on the calculation of the persistent current.
+Data Availability Statement
+The data that support the findings of this study are available from the corresponding author upon reasonable request. *
+Appendix A Results for higher generation system
+We remark that, as our model demand a very complicated implementation scheme in an actual real-life experimental situation, it could be difficult to fabricate and do the experiment for a large system.
+However, to convince the readers, in this appendix, we show the results for a 5th generation system (Fig.
+8) with $\mathcal{N}(5)=366$ sites as displayed in Fig.
+9.
+Figure 8: Schematic representation of the $5$th generation SPG fractal lattice geometry with hierarchical distribution of magnetic fluxes in the downward triangular plaquettes.
+Different colors represent the hierarchical flux distribution according to the order of hierarchy of the downward triangles (from smallest to the largest), viz., $\Phi_{1}$, $\Phi_{2}$, $\Phi_{3}$, $\Phi_{4}$, and $\Phi_{5}$.
+Figure 9: Energy eigenvalue spectrum for a 5th generation SPG lattice structure (having $\mathcal{N}(5)=366$ sites) as a function of the hierarchical magnetic flux with values of the hierarchy parameter chosen to be same as in Fig.
+2 and Fig.
+3.
+We found that, all our predictions from the previous generation systems (discussed in the main text) holds good for this relatively bigger system size also.
+It is worth mentioning that, as we have a deterministic disorder distribution of AB phases along different bonds for any given generation for our lattice model, there is no ready-made algorithm to generate the Hamiltonian for the system – one has do it by hand for each given generation.
+Figure 10: Schematic representation of a SPG fractal structure with a uniform distribution of the magnetic flux in all the upward triangular loops (left panel) and its energy eigenvalue distribution (right panel).
+In this study, we have considered the distribution of the magnetic flux in the downward triangular loops of the SPG fractal structure in such a way that it follows a special kind of hierarchical pattern.
+However, it is also interesting see what will happen if we put the magnetic flux in all the upward triangular loops of the SPG structure.
+This will impose a uniform Aharonov-Bohm phase factor along all the bonds in the SPG fractal structure (see Fig.
+10).
+In this case, there will be no concept of hierarchical flux distribution as automatically imposed by the ramification pattern of the SPG fractal geometry.
+We have done the computation for this case also and the results for the energy eigenvalue spectrum for a finite generation SPG fractal structure is shown in Fig.
+10.